OpenStudy (anonymous):

what is the value of the discriminate in the equation 9x^2-6x-8=0

OpenStudy (precal):

b^2-4ac is the discriminate

OpenStudy (anonymous):

so i square the b coefficient first?

OpenStudy (anonymous):

b^2-4*a*c =(-6)^2 - 4*9*(-8) =36+288 =324